# Adaptive Federated Learning with Auto-Tuned Clients

### v1

@@ 3 DELTA($\Delta$)-SGD: Distributed locality adaptive SGD
In Malitsky and Mishchenko (2020), a novel step size rule for (centralized) GD was introduced, which adapts to the minimization function’s local geometry. The proposed step size is clever and strikingly simple:
$\eta_{t}=\min\Big{\{}\tfrac{\|x_{t}-x_{t-1}\|}{2\|\nabla f(x_{t})-\nabla f(x_{t-1})\|},\sqrt{1+\theta_{t-1}}\eta_{t-1}\Big{\}},\quad\theta_{t-1}=\eta_{t-1}/\eta_{t-2}.$ (4)
The step size in (4) adapts to the local smoothness of the iterates, that is:
$\|\nabla f(x_{t})-\nabla f(x_{t-1})\|\leqslant L_{t}\cdot\|x_{t}-x_{t-1}\|,\quad\forall t=1,2,\dots$ (5)
and hence can increase during iterations, while the second condition in (4) ensures that $\eta_{t}$ does not increase arbitrarily. By definition, $L_{t}$ in (5) is never bigger than the global $L$ in (2), resulting in faster convergence in practice (Malitsky and Mishchenko, 2020, Section 4). Note that (4) is vastly different from some standard adaptive methods like Adagrad Duchi et al. (2011), which can only decrease the learning rate, as they accumulate non-negative quantities in the denominator of the step size. As a result, the scalar multiplied to their “adaptive” learning rates play a crucial role and requires proper tuning, similarly to vanilla SGD, as evident in Figure 1; see also Table 1 in Section 4.
We now introduce DELTA($\Delta$)-SGD: DistributEd LocaliTy Adaptive SGD, where we extend (4) to the FL setting, with the inclusion of stochasticity and local iterations, as summarized in Algorithm 1. For brevity, we use ~$\tilde{\nabla}f_{i}(x)=\frac{1}{|\mathcal{B}|}\sum\limits_{z\in\mathcal{B}}F_{i}(x,z)$ to denote the stochastic gradients with batch size $|\mathcal{B}|=b.$
@@
10:              $\theta_{t,k}^{i}=\eta_{t,k}^{i}/\eta_{t,k-1}^{i}$triangleright (line 10 & 11) update locality adaptive step size
@@
13:     $x_{t+1}=\frac{1}{|\mathcal{S}_{t}|}\sum\limits_{i\in\mathcal{S}_{t}}x_{t,K}^{i}$triangleright server-side aggregation
@@
We make a few remarks about Algorithm 1. First, the input $\eta_{0}>0$ can be quite arbitrary, as it can be corrected, per client level, in the 1st local iteration (line 9); similarly for $\theta_{0}>0$ (line 10). Second, as the step size $\eta_{t,k}^{i}$ requires two successive stochastic gradients, there could be two options: using $i)$ the same batch or $ii)$ a new batch; the second option will incur additional memory requirement, so we use the first. [^1] This distinction was already analyzed in the centralized setting, and it was reported that using the same batch performed better in practice Malitsky and Mishchenko (2020). Third, we include the “amplifier” $\gamma$ to the first condition of step size (line 9), but this is only needed for Theorem 1.222In practice, we did not change from the default value $\gamma=2$ in the original implementation in https://github.com/ymalitsky/adaptive_GD/blob/master/pytorch/optimizer.py. fourth, after local steps are done, $\eta_{0}$ is re-used (line 6); we could potentially use a more sophisticated approach and use a similar adaptive step size to line 9, but we leave this for future work. Last, we present simple averaging as the server-side aggregation for simplicity (line 13), but more sophisticated methods like FedAdam Reddi et al. (2021) can be used instead.
@@ 3.1 Convergence Analysis
Technical challenges. Before we present the convergence analysis, let us elaborate on the non-trivial technical challenges of analyzing Algorithm 1. The main difference of analyzing Algorithm 1 compared to other distributed/decentralized optimization theory is that the step size $\eta_{t,k}^{i}$ not only depends on the round $t$ and local steps $k$ (which is a challenge shared by other analyses), but also on $i$, the client. To deal with the client-dependent step size, we require a slightly non-standard assumption on the dissimilarity between $f$ and $f_{i}$, as shown below.
@@
$\mathbb{E}\|\nabla F_{i}(x,z)-\nabla f_{i}(x)\|$ $\leqslant\sigma^{2},$ (bounded variance) (1a)
@@
Assumption 1a has been standard in stochastic optimization literature Moulines and Bach (2011); Ghadimi and Lan (2013); Stich (2018); Khaled et al. (2020); recently, this assumption has been relaxed to weaker ones such as weak growth assumption Bottou et al. (2018) or expected smoothness Gower et al. (2021), but this is out of scope of this work. Assumption 1b is fairly standard in nonconvex optimization Zaheer et al. (2018); Ward et al. (2020), and often used in FL setting Xie et al. (2019, 2019); Reddi et al. (2021). Assumption 1c is reminiscent of the strong growth assumption in stochastic optimization literature Schmidt and Roux (2013), which is still used in recent works Cevher and Vũ (2019); Vaswani et al. (2019); Vaswani et al. (2019). To the best of our knowledge, this is the first theoretical analysis of the FL setting where clients can use their step size. We now present the main convergence theorem of Algorithm 1.
@@
where $\Psi_{1}=\max\left\{\frac{\sigma^{2}}{b},f(x_{0})-f(x^{\star})\right\}$ and $\Psi_{2}=\left(\frac{\sigma^{2}}{b}+G^{2}\right)$ are global constants, with $b=|\mathcal{B}|$ being the batch size; $\tilde{L}$ is a constant at most the average of local smoothness, i.e., $\max_{t}\frac{1}{m}\sum_{i=1}^{m}L_{i,t},$ where $L_{i,t}$ the local smoothness of $f_{i}$ at round $t$.
The convergence result in Theorem 1 implies a sublinear convergence to an $\varepsilon$-first order stationary point with at least $T=\mathcal{O}(\varepsilon^{-2})$ communication rounds. We emphasize again that the conditions on $\gamma$ and $\eta_{0}$ are only required for our theory. [^3] In all experiments in Section 4, we use the default setting $\gamma=2$, and $\eta_{0}=0.2$, without additional tuning. Importantly, we did not assume $f_{i}$ is $L$-smooth (or even $L_{i}$-smooth), a standard assumption in nonconvex optimization and FL literature Ward et al. (2020); Koloskova et al. (2020); Li et al. (2020); Reddi et al. (2021); instead, we can obtain a smaller quantity, $\tilde{L}$, through our analysis in a similar vein to (5); for space limitation, we defer the details and the proof to Appendix A.
@@ 4 Experimental Setup and Results
Datasets and models. To test the performance of different client optimizers in various settings, we evaluate on four datasets commonly used in FL scenarios: MNIST, FMNIST, CIFAR-10, and CIFAR-100 Krizhevsky et al. (2009). For MNIST and FMNIST, we train a shallow CNN with two convolutional and two fully-connected layers, followed by dropout and ReLU activations. For CIFAR-10, we train a ResNet-18 He et al. (2016). For CIFAR-100, we train both a ResNet-18 and a ResNet-50 to study the effect of changing the model architecture.
@@
For each dataset, we create a federated version by randomly partitioning the training data among $100$ clients, with each client getting $500$ examples. [^4] All the datasets we consider have $50,000$ training samples, leading to $500$ samples distributed per client. To control the level of “non-iidness”, we apply latent Dirichlet allocation (LDA) over the labels following Hsu et al. (2019), where each client has an associated multinomial distribution over the labels from which its examples are drawn. That is, the local training examples of each client has been drawn from a categorical distribution over $N$ classes, parameterized by $\mathbf{q}$, such that $\mathbf{q}\sim\text{Dir}(\alpha\mathbf{p}),$ where $\mathbf{p}$ is a prior distribution over $N$ classes, with $\alpha>0$ being the concentration parameter. We vary $\alpha\in\{0.01,0.1,1\}$, where bigger $\alpha$ indicates settings closer to i.i.d. scenarios. The class distribution of different $\alpha$’s is visualized in Figure 5 in Appendix B.
@@
FL setup and optimizers. For all cases, we fix the number of clients to be $100,$ and randomly sample $10\%$ among the $100$ clients. Similarly to Reddi et al. (2021); Li et al. (2020), we perform $E$ local epochs of training over each client’s dataset, and we utilize mini-batch gradients of size $b=64$, leading to $K\approx\lfloor\tfrac{E\cdot 500}{64}\rfloor$ local gradient steps; we use $E=1$ for all settings. For client optimizers, we compare stochastic gradient descent (SGD), SGD with momentum (SGDM), adaptive methods including Adam Kingma and Ba (2014), Adagrad Duchi et al. (2011), SGD with stochastic Polyak step size (SPS) Loizou et al. (2021), and our proposed method: $\Delta$-SGD in Algorithm 1. As our focus is on client adaptivity, we only present the results using simple FedAvg McMahan et al. (2017) as the server optimizer.
[…]
Implementation. We use Pytorch Paszke et al. (2019) to implement all experiments and run on eight P100 GPUs. Our implementation will be available online after the anonymous review process for reproducibility.
Figure 2: The effect of stronger heterogeneity on different client optimizers, induced by the Dirichlet concentration parameter $\alpha$.
@@ 4.1 Results
Results of the experiments detailed in the previous subsection are presented in Table 1. Below, we make comprehensive remarks, depending on: $i)$ the level of heterogeneity induced by different $\alpha$’s of the Dirichlet concentration parameter, $ii)$ the effect of changing the model architecture, and $iii)$ the effect of changing the dataset, and how these affect the performance of different client optimizers.
@@
$\alpha=1$ SGD 98.3${}_{\downarrow(0.2)}$ 86.5${}_{\downarrow(0.8)}$ 87.7${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(2.1)}}$ 57.7${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(4.2)}}$ 53.0${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(12.8)}}$
SGD ($\downarrow$) 97.8${}_{\downarrow(0.7)}$ 86.3${}_{\downarrow(1.0)}$ 87.8${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(2.0)}}$ 61.9${}_{\downarrow(0.0)}$ 60.9${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(4.9)}}$
SGDM 98.5${}_{\downarrow(0.0)}$ 85.2${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(2.1)}}$ 88.7${}_{\downarrow(1.1)}$ 58.8${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(3.1)}}$ 60.5${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(5.3)}}$
SGDM ($\downarrow$) 98.4${}_{\downarrow(0.1)}$ 87.2${}_{\downarrow(0.1)}$ 89.3${}_{\downarrow(0.5)}$ 61.4${}_{\downarrow(0.5)}$ 63.3${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(2.5)}}$
Adam 94.7${}_{{\color[rgb]{0.96,0.29,0.54}{\color[rgb]{0.96,0.29,0.54}\downarrow(3.8)}}}$ 71.8${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(15.5)}}$ 89.4${}_{\downarrow(0.4)}$ 55.6${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(6.3)}}$ 61.4${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(4.4)}}$
Adagrad 64.3${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(34.2)}}$ 45.5${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(41.8)}}$ 86.6${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(3.2)}}$ 53.5${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(8.4)}}$ 51.9${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(13.9)}}$
SPS 10.1${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(88.4)}}$ 85.9${}_{\downarrow(1.4)}$ 82.7${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(7.1)}}$ 1.0${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(60.9)}}$ 50.0${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(15.8)}}$
$\Delta$-SGD 98.4${}_{\downarrow(0.1)}$ 87.3${}_{\downarrow(0.0)}$ 89.8${}_{\downarrow(0.0)}$ 61.5${}_{\downarrow(0.4)}$ 65.8${}_{\downarrow(0.0)}$
$\alpha=0.1$ SGD 98.1${}_{\downarrow(0.0)}$ 83.6${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(2.8)}}$ 72.1${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(12.9)}}$ 54.4${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(6.7)}}$ 44.2${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(19.9)}}$
SGD ($\downarrow$) 98.0${}_{\downarrow(0.1)}$ 84.7${}_{\downarrow(1.7)}$ 78.4${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(6.6)}}$ 59.3${}_{\downarrow(1.8)}$ 48.7${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(15.4)}}$
SGDM 97.6${}_{\downarrow(0.5)}$ 83.6${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(2.8)}}$ 79.6${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(5.4)}}$ 58.8${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(2.3)}}$ 52.3${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(11.8)}}$
SGDM ($\downarrow$) 98.0${}_{\downarrow(0.1)}$ 86.1${}_{\downarrow(0.3)}$ 77.9${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(7.1)}}$ 60.4${}_{\downarrow(0.7)}$ 52.8${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(11.3)}}$
Adam 96.4${}_{\downarrow(1.7)}$ 80.4${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(6.0)}}$ 85.0${}_{\downarrow(0.0)}$ 55.4${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(5.7)}}$ 58.2${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(5.9)}}$
Adagrad 89.9${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(8.2)}}$ 46.3${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(40.1)}}$ 84.1${}_{\downarrow(0.9)}$ 49.6${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(11.5)}}$ 48.0${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(16.1)}}$
SPS 96.0${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(2.1)}}$ 85.0${}_{\downarrow(1.4)}$ 70.3${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(14.7)}}$ 42.2${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(18.9)}}$ 42.2${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(21.9)}}$
$\Delta$-SGD 98.1${}_{\downarrow(0.0)}$ 86.4${}_{\downarrow(0.0)}$ 84.5${}_{\downarrow(0.5)}$ 61.1${}_{\downarrow(0.0)}$ 64.1${}_{\downarrow(0.0)}$
$\alpha=0.01$ SGD 96.8${}_{\downarrow(0.7)}$ 79.0${}_{\downarrow(1.2)}$ 22.6${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(11.3)}}$ 30.5${}_{\downarrow(1.3)}$ 24.3${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(7.1)}}$
SGD ($\downarrow$) 97.2${}_{\downarrow(0.3)}$ 79.3${}_{\downarrow(0.9)}$ 33.9${}_{\downarrow(0.0)}$ 30.3${}_{\downarrow(1.5)}$ 24.6${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(6.8)}}$
SGDM 77.9${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(19.6)}}$ 75.7${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(4.5)}}$ 28.4${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(5.5)}}$ 24.8${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(7.0)}}$ 22.0${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(9.4)}}$
SGDM ($\downarrow$) 94.0${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(3.5)}}$ 79.5${}_{\downarrow(0.7)}$ 29.0${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(4.9)}}$ 20.9${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(10.9)}}$ 14.7${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(16.7)}}$
Adam 80.8${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(16.7)}}$ 60.6${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(19.6)}}$ 22.1${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(11.8)}}$ 18.2${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(13.6)}}$ 22.6${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(8.8)}}$
Adagrad 72.4${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(25.1)}}$ 45.9${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(34.3)}}$ 12.5${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(21.4)}}$ 25.8${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(6.0)}}$ 22.2${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(9.2)}}$
SPS 69.7${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(27.8)}}$ 44.0${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(36.2)}}$ 21.5${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(12.4)}}$ 22.0${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(9.8)}}$ 17.4${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(14.0)}}$
$\Delta$-SGD 97.5${}_{\downarrow(0.0)}$ 80.2${}_{\downarrow(0.0)}$ 31.6${}_{{\color[rgb]{0.96,0.29,0.54}\downarrow(2.3)}}$ 31.8${}_{\downarrow(0.0)}$ 31.4${}_{\downarrow(0.0)}$

### v3

@@ 3 DELTA($\Delta$)-SGD: Distributed locality adaptive SGD
We now introduce $\Delta$-SGD. In its simplest form, client $i$ at communication round $t$ uses the step size:
$\eta_{t}^{i}=\min\Big{\{}\tfrac{\|x_{t}^{i}-x_{t-1}^{i}\|}{2\|\nabla f_{i}(x_{t}^{i})-\nabla f_{i}(x_{t-1}^{i})\|},\sqrt{1+\theta_{t-1}^{i}}\eta_{t-1}^{i}\Big{\}},\quad\theta_{t-1}^{i}=\eta_{t-1}^{i}/\eta_{t-2}^{i}.$ (4)
The first part of $\min\{\cdot,\cdot\}$ approximates the (inverse of) local smoothness of $f_{i}$, [^1] Notice that $\|\nabla f_{i}(x_{t}^{i})-\nabla f_{i}(x_{t-1}^{i})\|\leqslant\tfrac{1}{2\eta_{t}^{i}}\|x_{t}^{i}-x_{t-1}^{i}\|\approx\tilde{L}_{i,t}\|x_{t}^{i}-x_{t-1}^{i}\|.$ and the second part controls how fast $\eta_{t}^{i}$ can increase. Indeed, $\Delta$-SGD with (4) enjoys the following decrease in the Lyapunov function:
$\|x_{t+1}-x^{\star}\|^{2}+\frac{1}{2m}\sum_{i=1}^{m}\|x_{t+1}^{i}-x_{t}^{i}\|^{2}+\frac{2}{m}\sum_{i=1}^{m}\Big{[}\eta_{t+1}^{i}\theta_{t+1}^{i}\big{(}f_{i}(x_{t}^{i})-f_{i}(x^{\star})\big{)}\Big{]}$
$\qquad\leqslant\|x_{t}-x^{\star}\|^{2}+\frac{1}{2m}\sum_{i=1}^{m}\|x_{t}^{i}-x_{t-1}^{i}\|^{2}+\frac{2}{m}\sum_{i=1}^{m}\Big{[}\eta_{t}^{i}\theta_{t}^{i}\big{(}f_{i}(x_{t-1}^{i})-f_{i}(x^{\star})\big{)}\Big{]},$ (5)
when $f_{i}$’s are assumed to be convex (c.f., Theorem 5 in Appendix A). For the FL settings, we extend (4) by including stochasticity and local iterations, as summarized in Algorithm 1 and visualized in Figure 7. For brevity, we use ~$\tilde{\nabla}f_{i}(x)=\frac{1}{|\mathcal{B}|}\sum_{z\in\mathcal{B}}\nabla F_{i}(x,z)$ to denote the stochastic gradients with batch size $|\mathcal{B}|=b$.
@@
10:              $\theta_{t,k}^{i}=\eta_{t,k}^{i}/\eta_{t,k-1}^{i}$triangleright (line 9 & 10) update locality adaptive step size
@@
13:     $x_{t+1}=\frac{1}{|\mathcal{S}_{t}|}\sum_{i\in\mathcal{S}_{t}}x_{t,K}^{i}$triangleright server-side aggregation
@@
We make a few remarks of Algorithm 1. First, the input $\theta_{0}>0$ can be quite arbitrary, as it can be corrected, per client level, in the first local iteration (line 10); similarly for $\eta_{0}>0$, although $\eta_{0}$ should be sufficiently small to prevent divergence in the first local step. Second, we include the “amplifier” $\gamma$ to the first condition of step size (line 9), but this is only needed for Theorem 1.222For all our experiments, we use the default value $\gamma=2$ from the original implementation in https://github.com/ymalitsky/adaptive_GD/blob/master/pytorch/optimizer.py. Last, ~$\tilde{\nabla}f_{i}(x_{t,k-1}^{i})$ shows up twice: in updating $x_{t,k}^{i}$ (line 8) and $\eta_{t,k}^{i}$ (line 9). Thus, one can use the same or different batches; we use the same batches in experiments to prevent additional gradient evaluations.
@@ 3.1 Convergence Analysis
Technical challenges. The main difference of analyzing Algorithm 1 compared to other decentralized optimization algorithms is that the step size $\eta_{t,k}^{i}$ not only depends on the round $t$ and local steps $k$, but also on $i$, the client. To deal with the client-dependent step size, we require a slightly non-standard assumption on the dissimilarity between $f$ and $f_{i}$, as detailed below.
@@
$\mathbb{E}\|\nabla F_{i}(x,z)-\nabla f_{i}(x)\|^{2}$ $\leqslant\sigma^{2},$ (bounded variance) (1a)
@@
Assumption 1a has been standard in stochastic optimization literature (Ghadimi & Lan, 2013; Stich, 2018; Khaled et al., 2020); recently, this assumption has been relaxed to weaker ones such as expected smoothness (Gower et al., 2021), but this is out of scope of this work. Assumption 1b is fairly standard in nonconvex optimization (Zaheer et al., 2018; Ward et al., 2020), and often used in FL setting (Xie et al., 2019a; b; Reddi et al., 2021). Assumption 1c is reminiscent of the strong growth assumption in stochastic optimization (Schmidt & Roux, 2013), which is still used in recent works (Cevher & Vũ, 2019; Vaswani et al., 2019a; b). To the best of our knowledge, this is the first theoretical analysis of the FL setting where clients can use their own step size. We now present the main theorem.
@@
where $\Psi_{1}=\max\left\{\frac{\sigma^{2}}{b},f(x_{0})-f(x^{\star})\right\}$ and $\Psi_{2}=\left(\frac{\sigma^{2}}{b}+G^{2}\right)$ are global constants, with $b=|\mathcal{B}|$ being the batch size; $\tilde{L}$ is a constant at most the maximum of local smoothness, i.e., $\max_{i,t}\tilde{L}_{i,t},$ where $\tilde{L}_{i,t}$ the local smoothness of $f_{i}$ at round $t$.
The convergence result in Theorem 1 implies a sublinear convergence to an $\varepsilon$-first order stationary point with at least $T=\mathcal{O}(\varepsilon^{-2})$ communication rounds. We remind again that the conditions on $\gamma$ and $\eta_{0}$ are only required for our theory. [^3] In all experiments, we use the default settings $\gamma=2$, $\eta_{0}=0.2$, and $\theta_{0}=1$ without additional tuning. Importantly, we did not assume $f_{i}$ is globally $L$-smooth, a standard assumption in nonconvex optimization and FL literature (Ward et al., 2020; Koloskova et al., 2020; Li et al., 2020; Reddi et al., 2021); instead, we can obtain a smaller quantity, $\tilde{L}$, through our analysis; for space limitation, we defer the details and the proof to Appendix A.
@@ 4 Experimental Setup and Results
We now introduce the experimental setup and discuss the results. Our implementation can be found in https://github.com/jlylekim/auto-tuned-FL.
@@
Datasets and models. We consider image classification and text classification tasks. We use four datasets for image classification: MNIST, FMNIST, CIFAR-10, and CIFAR-100 (Krizhevsky et al., 2009). For MNIST and FMNIST, we train a shallow CNN with two convolutional and two fully-connected layers, followed by dropout and ReLU activations. For CIFAR-10, we train a ResNet-18 (He et al., 2016). For CIFAR-100, we train both ResNet-18 and ResNet-50 to study the effect of changing the model architecture. For text classification, we use two datasets: DBpedia and AGnews datasets (Zhang et al., 2015), and train a DistillBERT (Sanh et al., 2019) for classification.
@@
For each dataset, we create a federated version by randomly partitioning the training data among $100$ clients ($50$ for text classification), with each client getting $500$ examples. To control the level of non-iidness, we apply latent Dirichlet allocation (LDA) over the labels following Hsu et al. (2019), where the degree class heterogeneity can be parameterized by the Dirichlet concentration parameter $\alpha.$ The class distribution of different $\alpha$’s is visualized in Figure 8 in Appendix B.9.
@@
FL setup and optimizers. We fix the number of clients to be $100$ for the image classification, and $50$ for the text classification; we randomly sample $10\%$ as participating clients. We perform $E$ local epochs of training over each client’s dataset. We utilize mini-batch gradients of size $b=64$ for the image classification, and $b=16$ for the text classification, leading to $K\approx\lfloor\tfrac{E\cdot 500}{b}\rfloor$ local gradient steps; we use $E=1$ for all settings. For client optimizers, we compare stochastic gradient descent (SGD), SGD with momentum (SGDM), adaptive methods including Adam (Kingma & Ba, 2014), Adagrad (Duchi et al., 2011), SGD with stochastic Polyak step size (SPS) (Loizou et al., 2021), and our proposed method: $\Delta$-SGD in Algorithm 1. As our focus is on client adaptivity, we mostly present the results using FedAvg (McMahan et al., 2017) as the server optimizer; additional results using FedAdam can be found in Appendix B.4.
Hyperparameters. For each optimizer, we perform a grid search of learning rates on a single task: CIFAR-10 classification trained with a ResNet-18, with Dirichlet concentration parameter $\alpha=0.1$; for the rest of the settings, we use the same learning rates. For SGD, we perform a grid search with $\eta\in\{0.01,0.05,0.1,0.5\}$. For SGDM, we use the same grid for $\eta$ and use momentum parameter $\beta=0.9$. To properly account for the SGD(M) fine-tuning typically done in practice, we also test dividing the step size by 10 after $50\%$, and again by 10 after $75\%$ of the total training rounds (LR decay). For Adam and Adagrad, we grid search with $\eta\in\{0.001,0.01,0.1\}.$ For SPS, we use the default setting of the official implementation. [^4] I.e., we use $f_{i}^{\star}=0$, and $c=0.5,$ for the SPS step size: $\frac{f_{i}(x)-f_{i}^{\star}}{c\|\nabla f_{i}(x)\|^{2}}$. The official implementation can be found in https://github.com/IssamLaradji/sps. For $\Delta$-SGD, we append $\delta$ in front of the second condition: $\sqrt{1+\delta\theta_{t,k-1}^{i}}\eta_{t,k-1}^{i}$ following Malitsky & Mishchenko (2020), and use $\delta=0.1$ for all experiments. [^5] We also demonstrate in Appendix B.1 that $\delta$ has very little impact on the final accuracy of $\Delta$-SGD. Finally, for the number of rounds $T$, we use $500$ for MNIST, $1000$ for FMNIST, $2000$ for CIFAR-10 and CIFAR-100, and $100$ for the text classification tasks.
@@ 4.1 Results
We clarify again that the best step size for each client optimizer was tuned via grid-search for a single task: CIFAR-10 classification trained with a ResNet-18, with Dirichlet concentration parameter $\alpha=0.1$. We then intentionally use the same step size in all other tasks to highlight two points: $i)$ $\Delta$-SGD works well without any tuning across different datasets, model architectures, and degrees of heterogeneity; $ii)$ other optimizers perform suboptimally without additional tuning.
@@
$\alpha=1$ SGD 98.3↓(0.2) 86.5↓(0.8) 87.7↓(2.1) 57.7↓(4.2) 53.0↓(12.8)
SGD ($\downarrow$) 97.8↓(0.7) 86.3↓(1.0) 87.8↓(2.0) 61.9↓(0.0) 60.9↓(4.9)
SGDM 98.5↓(0.0) 85.2↓(2.1) 88.7↓(1.1) 58.8↓(3.1) 60.5↓(5.3)
SGDM ($\downarrow$) 98.4↓(0.1) 87.2↓(0.1) 89.3↓(0.5) 61.4↓(0.5) 63.3↓(2.5)
Adam 94.7↓(3.8) 71.8↓(15.5) 89.4↓(0.4) 55.6↓(6.3) 61.4↓(4.4)
Adagrad 64.3↓(34.2) 45.5↓(41.8) 86.6↓(3.2) 53.5↓(8.4) 51.9↓(13.9)
SPS 10.1↓(88.4) 85.9↓(1.4) 82.7↓(7.1) 1.0↓(60.9) 50.0↓(15.8)
$\Delta$-SGD 98.4↓(0.1) 87.3↓(0.0) 89.8↓(0.0) 61.5↓(0.4) 65.8↓(0.0)
$\alpha=0.1$ SGD 98.1↓(0.0) 83.6↓(2.8) 72.1↓(12.9) 54.4↓(6.7) 44.2↓(19.9)
SGD ($\downarrow$) 98.0↓(0.1) 84.7↓(1.7) 78.4↓(6.6) 59.3↓(1.8) 48.7↓(15.4)
SGDM 97.6↓(0.5) 83.6↓(2.8) 79.6↓(5.4) 58.8↓(2.3) 52.3↓(11.8)
SGDM ($\downarrow$) 98.0↓(0.1) 86.1↓(0.3) 77.9↓(7.1) 60.4↓(0.7) 52.8↓(11.3)
Adam 96.4↓(1.7) 80.4↓(6.0) 85.0↓(0.0) 55.4↓(5.7) 58.2↓(5.9)
Adagrad 89.9↓(8.2) 46.3↓(40.1) 84.1↓(0.9) 49.6↓(11.5) 48.0↓(16.1)
SPS 96.0↓(2.1) 85.0↓(1.4) 70.3↓(14.7) 42.2↓(18.9) 42.2↓(21.9)
$\Delta$-SGD 98.1↓(0.0) 86.4↓(0.0) 84.5↓(0.5) 61.1↓(0.0) 64.1↓(0.0)
$\alpha=0.01$ SGD 96.8↓(0.7) 79.0↓(1.2) 22.6↓(11.3) 30.5↓(1.3) 24.3↓(7.1)
SGD ($\downarrow$) 97.2↓(0.3) 79.3↓(0.9) 33.9↓(0.0) 30.3↓(1.5) 24.6↓(6.8)
SGDM 77.9↓(19.6) 75.7↓(4.5) 28.4↓(5.5) 24.8↓(7.0) 22.0↓(9.4)
SGDM ($\downarrow$) 94.0↓(3.5) 79.5↓(0.7) 29.0↓(4.9) 20.9↓(10.9) 14.7↓(16.7)
Adam 80.8↓(16.7) 60.6↓(19.6) 22.1↓(11.8) 18.2↓(13.6) 22.6↓(8.8)
Adagrad 72.4↓(25.1) 45.9↓(34.3) 12.5↓(21.4) 25.8↓(6.0) 22.2↓(9.2)
SPS 69.7↓(27.8) 44.0↓(36.2) 21.5↓(12.4) 22.0↓(9.8) 17.4↓(14.0)
$\Delta$-SGD 97.5↓(0.0) 80.2↓(0.0) 31.6↓(2.3) 31.8↓(0.0) 31.4↓(0.0)